Debt payments of $400 due today and $900 due in three years are to be combined into a single payment due two years from now. What is the size of the single payment if interest is 8% compounded quarterly? ___________

Need it in 30 minutes


1
Expert's answer
2021-02-24T06:51:43-0500

For the first payment:

"FV_1=PV_1(1+i)^n"

where FV is future value, PV is present value, i is the interest per compounding period, n is the number of compounding periods.

"FV_1=400(1+0.08/4)^{4\cdot2}=\$468.66"

For the second payment:

"PV_2=FV_2(1+i)^{-n}"

"PV_2=900(1+0.02)^{-1\cdot4}=\$831.46"


Single payment:

"FV=FV_1+PV_2=468.66+831.46=\$1300.12"